OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and in the air with an initial velocity of 42 feet per second. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and in the air with an initial velocity of 42 feet per second. Approximately what maximum height does the soccer ball reach? I think this can not be solve due to lack of data.

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

Here's the for choices a 1.3 secs b 2.6 secs c 16.0 secs d 27.6 secs

OpenStudy (anonymous):

some options are available or not

OpenStudy (anonymous):

^^^ i put them

OpenStudy (anonymous):

kk

OpenStudy (lgbasallote):

use the formula to get maximum height first \[S_2 - S_1 = V_1 t + \frac12 at\] s2 is the maximum height = not given s1 is initial height = 0 V1 is the initial velocity = 42 a = -9.8 m/s^2 t = time =not given

OpenStudy (lgbasallote):

obviously that wont work ;)

OpenStudy (anonymous):

you have to deal with two components of the balls motion ie vertical and horizontal

OpenStudy (lgbasallote):

use another formula \[V_2 = V_1 + at\] v2 is final speed which will be 0 because v = 0 at max height v1 = 42 a = -9.8 now solve this... \[0 = 42 -9.8t\] \[9.8t = 42\] can you solve for t?

OpenStudy (anonymous):

No this is for my bro i dont know how to

OpenStudy (lgbasallote):

hmm if you look closely..that last step is just algebra...

OpenStudy (anonymous):

o ok thanks for the help guys

OpenStudy (lgbasallote):

btw...you need to multiply the time you get from that by 2 because the one we got was just the first half

OpenStudy (anonymous):

I got 8.6 and thats not one of the answers.